# Selmer stability for elliptic curves in Galois $\ell$-extensions

### v1

@@ 4. Density results and connections with Malle’s conjecture
  • the mod $\ell$ Galois representation $\rho_{E,\ell}:\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ is surjective.
Let $G$ be a finite $\ell$-group with #$\#G=\ell^{n}$. We recall that Theorem 3.10 establishes the existence of infinitely many Galois extensions $L/\mathbb{Q}$ such that $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ and $\operatorname{Sel}_{\ell}(E/L)=0$. In this section, we go further and obtain asymptotic lower bounds for the number of such extensions $L/\mathbb{Q}$, counted by their absolute discriminant. We then compare these lower bounds with the expected asymptotics for the total number of $G$-extensions of $\mathbb{Q}$, as predicted by the weak form of Malle’s conjecture.
@@ 4.1. Counting $\ell$-extensions by discriminant
$1\rightarrow\mathbb{Z}/\ell\mathbb{Z}\rightarrow G\rightarrow\overline{G}\rightarrow 1$ (4.1)
Let $L/\mathbb{Q}$ be a Galois extension with $\operatorname{Gal}(L/\mathbb{Q})\simeq\overline{G}$ and $\widetilde{L}/L/\mathbb{Q}$ a solution to the above embedding problem. Let $\zeta$ be a primitive $\ell$-th root of unity and let $\sigma$ be a generator of $\operatorname{Gal}(\widetilde{L}(\mu_{\ell})/\widetilde{L})$ with $\sigma(\zeta)=\zeta^{q}$.
@@
The assertions (i) and (ii) follow from Proposition 3.3 and 3.4 of [KM04] respectively. ∎
Let $\mathbb{Q}_{1}$ be the unique $\mathbb{Z}/\ell\mathbb{Z}$-extension of $\mathbb{Q}$ which is contained in $\mathbb{Q}(\mu_{\ell^{2}})$. Let $b\in\mathbb{Q}(\mu_{\ell})^{\times}$ be such that $\sigma(b)/b^{q}\in(\mathbb{Q}(\mu_{\ell})^{\times})^{\ell}$. Setting $L:=\mathbb{Q}$ and $\overline{G}=1$, we choose $\widetilde{L}:=\mathbb{Q}_{1}$. Then, any other cyclic $\ell$-extension of $\mathbb{Q}$ is obtained by twisting $\mathbb{Q}_{1}$ by an element $b$. One denotes this twist by $\mathbb{Q}_{b}$ and thus we have an indexing of all cyclic $\ell$-extensions of $\mathbb{Q}$.
@@
The result follows from [KM04, Proposition 3.5]. ∎
@@
$\operatorname{Hom}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})\simeq\{b\in\mathbb{Q}(\mu_{\ell})^{\times}/(\mathbb{Q}(\mu_{\ell})^{\times})^{\ell}:\sigma(b)/b^{q}\in(\mathbb{Q}(\mu_{\ell})^{\times})^{\ell}\}.$
@@ 4.2. An application of Wiles’ formula
Let $S$ be a finite set of primes in $\mathfrak{T}_{E,L}$ disjoint from $\Sigma\cup\{p_{1},\dots,p_{m}\}$. Let $V_{S}$ be the space of cohomology classes $f\in H^{1}(\mathbb{Q}_{S}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})$ such that $\operatorname{res}_{q}(f)=0$ for all primes $q\in\Sigma\cup\{p_{1},\dots,p_{m}\}$. Setting $\widetilde{S}:=S\cup\Sigma\cup\{p_{1},\dots,p_{m}\}$, the space $V_{S}=H^{1}_{\mathcal{L}}(\mathbb{Q}_{\widetilde{S}}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})$ is determined by Selmer conditions $\{\mathcal{L}_{q}\}_{q\in\widetilde{S}}$ defined as follows:
$\mathcal{L}_{q}:=\begin{cases}H^{1}(\mathbb{Q}_{q},\mathbb{Z}/\ell\mathbb{Z})&\text{ if }q\in S\\ 0&\text{ if }q\in\Sigma\cup\{p_{1},\dots,p_{m}\}=.\end{cases}$
Let $\overline{\chi}$ be the mod $\ell$ cyclotomic character. The dual Selmer group $V_{S}^{\perp}:=H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{\widetilde{S}}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z}(\overline{\chi}))$ is given by:
@@
Let $\mathbb{Q}_{f}$ be the cyclic $\ell$-extension of $\mathbb{Q}$ which is cut out by $f$ and set $\Delta_{f}:=\Delta_{\mathbb{Q}_{f}}$. We have that $|\Delta_{f}|\leq(\prod_{q\in S}q)^{\ell-1}$ and part (ii) follows from Proposition 4.2. ∎
@@
Let $\psi_{1},\dots,\psi_{r}$ be the non-zero elements in $H^{1}(\mathbb{Q}_{\Sigma\cup\{p_{1},\dots,p_{m}\}}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z}(\overline{\chi}))$. Let $\psi\in\{\psi_{1},\dots,\psi_{r}\}$ and $\Omega:=\operatorname{Gal}(\mathbb{Q}(\mu_{\ell})/\mathbb{Q})$. Note that the inflation–restriction sequence gives:
@@
Since #$\#\Omega$ is coprime to $\ell$, $H^{1}\left(\Omega,\mathbb{Z}/\ell\mathbb{Z}(\overline{\chi})\right)=0$ and thus the restriction map is an injection. Thus the homomorphism $\operatorname{res}(\psi)$ is non-zero. It gives rise to a $\mathbb{Z}/\ell\mathbb{Z}$ extension $L_{\psi}$ over $\mathbb{Q}(\mu_{\ell})$, which is a Galois extension of $\mathbb{Q}$ with
@@
There are two possibilities for the intersection $L_{\psi}\cap L(\mu_{\ell})$, namely either $L_{\psi}\cap L(\mu_{\ell})=L_{\psi}$ or $L_{\psi}\cap L(\mu_{\ell})=\mathbb{Q}(\mu_{\ell})$. Recall that $\Omega$ acts non-trivially on $\operatorname{Gal}(L_{\psi}/\mathbb{Q}(\mu_{\ell}))$. On the other hand, $\operatorname{Gal}(L(\mu_{\ell})/\mathbb{Q})\simeq\Omega\times\operatorname{Gal}(L/\mathbb{Q})$. Therefore, if $L_{\psi}\subset L(\mu_{\ell})$, then the action of $\Omega$ on $\operatorname{Gal}(L_{\psi}/\mathbb{Q}(\mu_{\ell}))$ would be trivial, which contradicts (4.2). Thus we have that $L_{\psi}\cap L(\mu_{\ell})=\mathbb{Q}(\mu_{\ell})$. Lemma 3.8 implies that $L_{\psi}\cap\mathbb{Q}(E[\ell])=\mathbb{Q}(\mu_{\ell})$. The Chebotarev conditions determining $\mathfrak{T}_{E,L}$ are thus independent of the extension $L_{\psi}/\mathbb{Q}(\mu_{\ell})$. There is thus a nonempty Chebotarev set contained in $\operatorname{Gal}(L_{\psi}\cdot L(E[\ell])/\mathbb{Q})$ determining a positive density set of primes $q\in\mathfrak{T}_{E,L}$ which are non-split in the extension $L_{\psi}/\mathbb{Q}(\mu_{\ell})$. Pick one such prime $q_{i}$ for each $\psi_{i}$ for $i=1,\dots,r$. Then setting $S_{0}:=\{q_{1},\dots,q_{r}\}$, we have that:
@@
For the rest of this article, we fix a choice of $S_{0}$ as in Lemma 4.4 and set $Z:=S_{0}\cup\Sigma\cup\{p_{1},\dots,p_{m}\}$.
@@
[…]
@@
#$\#W_{T}=\#V_{S_{0}}\times\#\mathcal{W}=\#V_{S_{0}}(\ell-1)^{\#T}.$
As $T$ ranges over finite subsets of $\mathfrak{T}_{E,L}\setminus Z$, the sets $W_{T}$ are mutually disjoint. To see this in detail, let $T$ and $T^{\prime}$ be distinct finite subsets of $\mathfrak{T}_{E,L}\setminus Z$. Without loss of generality, suppose that $T\not\subset T^{\prime}$. Then there exists a prime $q\in T\setminus T^{\prime}$. By construction, every element of $W_{T}$ is ramified at $q$, whereas every element of $W_{T^{\prime}}$ is unramified at $q$. It follows that $W_{T}\cap W_{T^{\prime}}=\emptyset$. Let $W$ denote the disjoint union $W=\bigsqcup_{T}W_{T}$ as $T$ ranges over finite subsets of $\mathfrak{T}_{E,L}\backslash Z$.
@@
$g(s):=\sum_{T}(\ell-1)^{\#T}\left(\prod_{q\in T}q\right)^{-s}=\prod_{q\in\mathfrak{T}_{E,L}\backslash Z}\left(1+(\ell-1)q^{-s}\right),$
where $T$ ranges over finite subsets of $\mathfrak{T}_{E,L}\backslash Z$. Write $g(s)=\sum_{n\geq 1}a_{n}n^{-s}$ where $a_{n}:=(\ell-1)^{r}$ if $n$ is a product of $r$ distinct primes in $\mathfrak{T}_{E,L}\backslash Z$, and $a_{n}:=0$ otherwise.
Arguing as in [Ser75, Theorem 2.4, p.5], we find that
@@
where $\alpha:=\mathfrak{d}(\mathfrak{T}_{E,L})$. It follows from Lemma 3.7 that
@@
where $h(s)$ is a non-zero holomorphic function in $\operatorname{Re}(s)\geq 1$. By Delange’s Tauberian theorem (cf. [Ten15, Theorem 7.28]),
@@
$\mathcal{M}(G,E,X)\geq\#\left\{f\in W:\prod_{q\in T}q\leq c_{3}X^{\frac{1}{\ell^{n-1}(\ell-1)}}\right\},$ (4.6)
@@
$c_{3}:=\left(c_{2}^{\frac{1}{\ell^{n-1}(\ell-1)}}\prod_{q\in S_{0}}q\right)^{-1}.$
@@
#$\#\left\{f\in W:\prod_{q\in T}q\leq c_{3}X^{\frac{1}{\ell^{n-1}(\ell-1)}}\right\}=\sum_{n\leq c_{3}X^{\frac{1}{\ell^{n-1}(\ell-1)}}}a_{n}$ (4.7)

### v2

@@ 4. Density results and connections with Malle’s conjecture
  • the mod $\ell$ Galois representation $\rho_{E,\ell}\mathrel{\mathop{\ordinarycolon}}\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\to\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ is surjective.
Let $G$ be a finite $\ell$-group with $G=\ell^{n}$. We recall that Theorem 3.10 establishes the existence of infinitely many Galois extensions $L/\mathbb{Q}$ such that $\operatorname{Gal}(L/\mathbb{Q})\simeq G$ and $\operatorname{Sel}_{\ell}(E/L)=0$. In this section, we go further and obtain asymptotic lower bounds for the number of such extensions $L/\mathbb{Q}$, counted by their absolute discriminant. We then compare these lower bounds with the expected asymptotics for the total number of $G$-extensions of $\mathbb{Q}$, as predicted by the weak form of Malle’s conjecture.
@@ 4.1. Counting $\ell$-extensions by discriminant
$1\rightarrow\mathbb{Z}/\ell\mathbb{Z}\rightarrow G\rightarrow\overline{G}\rightarrow 1.$ (4.1)
[…]
@@
The assertions (i) and (ii) follow from Proposition 3.3 and 3.4 of [klunersmalle] respectively. ∎
Let $\mathbb{Q}_{1}$ be the unique $\mathbb{Z}/\ell\mathbb{Z}$-extension of $\mathbb{Q}$ which is contained in $\mathbb{Q}(\mu_{\ell^{2}})$. Let $b\in\mathbb{Q}(\mu_{\ell})^{\times}$ be such that $\sigma(b)/b^{q}\in(\mathbb{Q}(\mu_{\ell})^{\times})^{\ell}$. Setting $L\mathrel{\mathop{\ordinarycolon}}$$\mathbb{Q}$ and $\overline{G}=1$, we choose $\widetilde{L}\mathrel{\mathop{\ordinarycolon}}=\mathbb{Q}_{1}$. Then, any other cyclic $\ell$-extension of $\mathbb{Q}$ is obtained by twisting $\mathbb{Q}_{1}$ by an element $b$. One denotes this twist by $\mathbb{Q}_{b}$ and thus we have an indexing of all cyclic $\ell$-extensions of $\mathbb{Q}$.
@@
The result follows from [klunersmalle, Proposition 3.5]. ∎
@@
$\operatorname{Hom}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})\simeq\{b\in\mathbb{Q}(\mu_{\ell})^{\times}/(\mathbb{Q}(\mu_{\ell})^{\times})^{\ell}\mathrel{\mathop{\ordinarycolon}}\sigma(b)/b^{q}\in(\mathbb{Q}(\mu_{\ell})^{\times})^{\ell}\}.$
@@ 4.2. An application of Wiles’ formula
[…]
$\mathcal{L}_{q}\mathrel{\mathop{\ordinarycolon}}=\begin{cases}H^{1}(\mathbb{Q}_{q},\mathbb{Z}/\ell\mathbb{Z})&\text{ if }q\in S\\ 0&\text{ if }q\in\Sigma\cup\{p_{1},\dots,p_{m}\}.\end{cases}$
Let $\overline{\chi}$ be the mod $\ell$ cyclotomic character. The dual Selmer group $V_{S}^{\perp}\mathrel{\mathop{\ordinarycolon}}=H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{\widetilde{S}}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z}(\overline{\chi}))$ is given by:
@@
Let $\mathbb{Q}_{f}$ be the cyclic $\ell$-extension of $\mathbb{Q}$ which is cut out by $f$ and set $\Delta_{f}\mathrel{\mathop{\ordinarycolon}}=\Delta_{\mathbb{Q}_{f}}$. We have that $|\Delta_{f}|\leq(\prod_{q\in S}q)^{\ell-1}$ and part (ii) follows from Proposition 4.2. ∎
@@
Let $\psi_{1},\dots,\psi_{r}$ be the non-zero elements in $H^{1}(\mathbb{Q}_{\Sigma\cup\{p_{1},\dots,p_{m}\}}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z}(\overline{\chi}))$. Let $\psi\in\{\psi_{1},\dots,\psi_{r}\}$ and $\Omega\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\mathbb{Q}(\mu_{\ell})/\mathbb{Q})$. Note that the inflation–restriction sequence gives:
@@
Since $\Omega$ is coprime to $\ell$, $H^{1}\left(\Omega,\mathbb{Z}/\ell\mathbb{Z}(\overline{\chi})\right)=0$ and thus the restriction map is an injection. Thus the homomorphism $\operatorname{res}(\psi)$ is non-zero. It gives rise to a $\mathbb{Z}/\ell\mathbb{Z}$ extension $L_{\psi}$ over $\mathbb{Q}(\mu_{\ell})$, which is a Galois extension of $\mathbb{Q}$ with
@@
There are two possibilities for the intersection $L_{\psi}\cap L(\mu_{\ell})$, namely either $L_{\psi}\cap L(\mu_{\ell})=L_{\psi}$ or $L_{\psi}\cap L(\mu_{\ell})=\mathbb{Q}(\mu_{\ell})$. Recall that $\Omega$ acts non-trivially on $\operatorname{Gal}(L_{\psi}/\mathbb{Q}(\mu_{\ell}))$. On the other hand, $\operatorname{Gal}(L(\mu_{\ell})/\mathbb{Q})\simeq\Omega\times\operatorname{Gal}(L/\mathbb{Q})$. Therefore, if $L_{\psi}\subset L(\mu_{\ell})$, then the action of $\Omega$ on $\operatorname{Gal}(L_{\psi}/\mathbb{Q}(\mu_{\ell}))$ would be trivial, which contradicts (4.2). Thus we have that $L_{\psi}\cap L(\mu_{\ell})=\mathbb{Q}(\mu_{\ell})$. Lemma 3.8 implies that $L_{\psi}\cap\mathbb{Q}(E[\ell])=\mathbb{Q}(\mu_{\ell})$. The Chebotarev conditions determining $\mathfrak{T}_{E,L}$ are thus independent of the extension $L_{\psi}/\mathbb{Q}(\mu_{\ell})$. There is thus a nonempty Chebotarev set contained in $\operatorname{Gal}(L_{\psi}\cdot L(E[\ell])/\mathbb{Q})$ determining a positive density set of primes $q\in\mathfrak{T}_{E,L}$ which are non-split in the extension $L_{\psi}/\mathbb{Q}(\mu_{\ell})$. Pick one such prime $q_{i}$ for each $\psi_{i}$ for $i=1,\dots,r$. Then setting $S_{0}\mathrel{\mathop{\ordinarycolon}}=\{q_{1},\dots,q_{r}\}$, we have that:
@@
For the rest of this article, we fix a choice of $S_{0}$ as in Lemma 4.4 and set $Z\mathrel{\mathop{\ordinarycolon}}=S_{0}\cup\Sigma\cup\{p_{1},\dots,p_{m}\}$.
@@
$0\rightarrow V_{S_{0}}\xrightarrow{\iota}V_{S_{0}\cup T}\xrightarrow{\pi}\bigoplus_{q\in T}\frac{H^{1}(\mathbb{Q}_{q},\mathbb{Z}/\ell\mathbb{Z})}{H^{1}_{\operatorname{nr}}(\mathbb{Q}_{q},\mathbb{Z}/\ell\mathbb{Z})}.$ (4.3)
@@
$W_{T}=\#V_{S_{0}}\times\#\mathcal{W}=\#V_{S_{0}}(\ell-1)^{\#T}.$
As $T$ ranges over finite subsets of $\mathfrak{T}_{E,L}\setminus Z$, the sets $W_{T}$ are mutually disjoint. To see this in detail, let $T$ and $T^{\prime}$ be distinct finite subsets of $\mathfrak{T}_{E,L}\setminus Z$. Without loss of generality, suppose that $T\not\subset T^{\prime}$. Then there exists a prime $q\in T\setminus T^{\prime}$. By construction, every element of $W_{T}$ is ramified at $q$, whereas every element of $W_{T^{\prime}}$ is unramified at $q$. It follows that $W_{T}\cap W_{T^{\prime}}=\emptyset$. Let $W$ denote the disjoint union $W=\bigsqcup_{T}W_{T}$ as $T$ ranges over finite subsets of $\Z\mathfrak{T}_{E,L}\backslash$ Z.
@@
[…]
where $T$ ranges over finite subsets of $\Z\mathfrak{T}_{E,L}\backslash$ Z. Write $g(s)=\sum_{n\geq 1}a_{n}n^{-s}$ where $a_{n}\mathrel{\mathop{\ordinarycolon}}=(\ell-1)^{r}$ if $n$ is a product of $r$ distinct primes in $\Z\mathfrak{T}_{E,L}\backslash$ Z, and $a_{n}\mathrel{\mathop{\ordinarycolon}}=0$ otherwise.
Arguing as in [serredivisibilite, Theorem 2.4, p.5], we find that
@@
where $\alpha\mathrel{\mathop{\ordinarycolon}}=\mathfrak{d}(\mathfrak{T}_{E,L})$. It follows from Lemma 3.7 that
@@
where $h(s)$ is a non-zero holomorphic function in $\operatorname{Re}(s)\geq 1$. By Delange’s Tauberian theorem (cf. [Tenenbaum, Theorem 7.28]),
@@
$\mathcal{M}(G,E,X)\geq\#\left\{f\in W\mathrel{\mathop{\ordinarycolon}}\prod_{q\in T}q\leq c_{3}X^{\frac{1}{\ell^{n-1}(\ell-1)}}\right\},$ (4.6)
@@
$c_{3}\mathrel{\mathop{\ordinarycolon}}=\left(c_{2}^{\frac{1}{\ell^{n-1}(\ell-1)}}\prod_{q\in S_{0}}q\right)^{-1}.$
@@
$\left\{f\in W\mathrel{\mathop{\ordinarycolon}}\prod_{q\in T}q\leq c_{3}X^{\frac{1}{\ell^{n-1}(\ell-1)}}\right\}=\sum_{n\leq c_{3}X^{\frac{1}{\ell^{n-1}(\ell-1)}}}a_{n}$ (4.7)